# Understanding Self-Regulated Learning Behavior Among High and Low Dropout Risk Students During CS1: Combining Trace Logs, Dropout Prediction and Self-Reports

### v1

@@ 3.2. The study
The study was implemented in two phases. The initial study entailed completing the course and using the available learning resources. During the study, the trace log data, task points, exam points, and final grades of the participants were collected from the virtual learning environment. After the course, an additional elective post-course questionnaire was administered to the original participants to gain additional self-report data. We followed the university’s guidelines for conducting studies involving human participants. Prior to participation, students were shown a full privacy statement on how their data were to be used in this study. Participating in both studies was fully elective, and participants could leave the study at any time. Participation was based on informed consent, and whether a student participated in the study carried no incentives or penalties. Altogether, $N=47$ students participated in the initial study. Out of them, 29 students also answered the additional post-course questionnaire.
@@ 3.2.1. Data collection
The main data of the study are the trace logs of the participating students. The initial logs were collected as raw action logs which contained timestamps of every action a student performed in the virtual learning environment during the 11 weeks of the course. The raw logs were processed into final trace logs by removing actions irrelevant to the course and by recoding the actions into context-relevant event codes. The codes relate to the various learning resources (e.g., course materials, lectures, supplementary materials, group sessions, one-on-one sessions) students use and the different tasks (introductory, basic, core, bonus, guru) students completed during CS1. Additionally, the events relating to different weekly task sets were differentiated, as student could, for instance, view and complete tasks from the current week or review and check answers to tasks from a previous week. Overall, 82 distinct event codes were used to describe student behavior during CS1. The description of all event codes and the specific approach for processing the raw logs is described in Appendix A.
In addition to trace log data, task completion and performance data was also collected for the duration of the whole course. The data included timestamps of all task submission attempts sent to the virtual learning environment and the correctness of each submission as determined by whether the student got full points from the task. Additionally, general performance data, such as weekly task points before the weekly deadline, task points after the deadline (as students could edit their answers based on the model answers), final exam points and final course grade were collected.
@@ 3.2.2. Data analysis
Figure 2. The procedure for analyzing trace logs used in the present paper. Analysis follows Winne and Marzouk (2019)’s characterization of SRL, whereby student’s learning strategy are described as a transition of learning tactics. The approach for identifying learning tactics from trace logs is based on Matcha et al. (2019).
Two analyses were conducted to address the research questions. First, process mining used to examine students’ weekly learning strategies with course materials. Weekly dropout risk was calculated using a pre-trained predictor based on task completion data. Low- and high-risk strategies were then clustered to identify different learning types. Second, student profiles derived from these strategies were clustered and evaluated using self-reports for validation. Figure 2 outlines the full analysis procedure, with the details provided next. All anonymized data and analysis code is available online [^1] URL available upon request from the first author..
Before the main analysis, students’ weekly behavior patterns were extracted from trace logs. Building on the characterization of student SRL behavior via learning tactics and learning strategies (see Section 2.1), we first identified students’ learning tactics of using the course materials, and then we mapped strategies for applying the tactics during each course week. Since the course follows a weekly structure with consistent learning goals and tasks, we mapped strategies for each week. A process mining and clustering method, adapted from prior work (Zhidkikh et al., 2023), was employed. This method, based on event-based SRL analysis by Matcha et al. (2019) and Jovanović et al. (2017), involved splitting trace logs into per-student study sessions using the time distance between the events. Study sessions represent consecutive active interactions with course materials, such as student reading or completing a task. If the time difference between two events was longer than 25 minutes, the first event was set as an end of the previous session and the second as the start of a new session. We chose a 25-minute cut-off based on the number of weekly tasks and the average time students were expected to use on each task. Single-event sessions were then excluded to remove logs of accidental virtual learning environment interactions. Each session was represented as a frequency vector $(\text{fE}_{1},\text{fE}_{2},\ldots,\text{fE}_{82})$, where $\text{fE}_{i}$ indicates the relative frequency of an event type in the session. These vectors were clustered using $k$-medoids with the $\ell_{1}$ metric (Schubert and Lenssen, 2022) to enhance robustness of the analysis in case the dataset was too limited in size. The clusters were then interpreted as distinct learning tactics employed by students in CS1. Finally, each student’s weekly strategy was constructed by remapping the study sessions into tactic types and generating First-Order Markov Models of weekly tactic use through process discovery (Berti et al., 2023).
@@
Additionally, the open-ended questionnaire answers were preprocessed for later analysis. The last two authors of this study individually coded initial answers into themes. The coding of the themes was not guided, but emerged from the questionnaire answers (cf. conventional content analysis in Hsieh and Shannon, 2005). Additionally, the two authors recorded whether a student had a self-reported positive, neutral or negative opinion of their learning behavior during the course. After individual coding, the two authors convened in a two-hour session to form a consensus on codings which were not initially agreed upon. The resulting codes of all students are thus jointly decided between the last two authors. The codes were used in subsequent analyses to compare students and characterize their behavior.
To answer RQ1 and RQ2, weekly learning strategies were paired with weekly dropout predictions to identify patterns in the behavior of low- and high-risk students. Since students may adjust their strategies weekly in response to environmental factors and prior performance (e.g., Pintrich, 2000), their dropout risk may fluctuate over time. We therefore assessed dropout risk weekly with a model developed in a prior study (Zhidkikh et al., 2024), which enabled the identification of “low dropout risk” and “high dropout risk” strategies. The used model was trained on task performance data from seven past iterations of the CS1 course, allowing prediction of weekly dropout based on task performance. Given the task performance data, the prediction model gives a dropout prediction $p_{\text{drop}}$ as a probability value from 0 (likely to pass) to 1 (likely to dropout). The learning strategies, modeled as First-Order Markov Models, were categorized into low and high dropout risk groups using the weekly predictions. These groups were then clustered using the Expectation Maximization algorithm. The cluster counts were selected by the first author using cluster validation indices (CVIs) as a guide. The first author then devised a dashboard, in which for each cluster median dropout probability, median grade of students, combined self-report codes, combined heuristic net (Weijters et al., 2006), per-strategy weekly performance data, per-strategy First-Order Markov Model, and per-strategy SHAP waterfall plot explaining the dropout prediction (Lundberg and Lee, 2017) were made available for the first two authors to help characterize the learning strategy types. To answer the research questions, the clusters were interpreted as strategy types by comparing them to each other using the aforementioned data. The interpretation was done by the first two authors through two hours of joint work, where the clusters were reviewed and the interpretations agreed upon.
Additionally, the strategy types and student self-reports were utilized to formulate students’ course activity profiles to further address RQ1 and RQ2, and also RQ3. First, weekly student strategies were mapped to the strategy types identified in the analysis of RQ1 and RQ2. Then, each student was assigned an activity profile represented as a frequency vector $(\text{fStrat}_{1},\text{fStrat}_{2},\ldots,\text{fStrat}_{N})$, where $\text{fStrat}_{i}$ represents the frequency of a student’s weekly strategies corresponding to a specific learning strategy type, and $N$ is the number of strategy types identified earlier. In other words, these profiles capture students’ course activities and their strategy adjustments. The use of frequencies simplified profiles, enhancing robustness of the analysis. Agglomerative clustering with the $\ell_{1}$ metric was applied to cluster students based on their activity profiles. The final cluster count was selected by the first author using CVIs and the resulting dendrogram. The resulting clusters were interpreted by the first two authors iteratively in up to three hours of joint work using activity profiles, mean dropout prediction, course performance, and coded self-reports, with the last data source particularly aiding cluster interpretation. To support comparison of performance data between clusters, Brunner-Munzel test for stochastic equality ($H_{0}:P(X<Y)=P(X>Y)$) was used as a robust non-parametric alternative to the Mann-Whitney U test, suitable for low number of observations and the presence of tied values (Neubert and Brunner, 2007). This paper adopts the reporting guidelines proposed by Karch (2021), including using stochastic superiority estimate $\hat{p}^{\prime\prime}$ for effect size. While the effect size estimates for Brunner-Munzel test do not have standardized significance categories, $\hat{p}^{\prime\prime}$ can be interpreted as the probability that random variable $X$ is lower than random variable $Y$ when splitting ties (i.e. observations of equal values) evenly (Karch, 2021).

### v2

@@ 3.2. The study
The study was implemented in two phases. The initial study entailed completing the course and using the available learning resources. During the study, the trace log data, task points, exam points, and final grades of the participants were collected from the virtual learning environment. After the course, an additional elective post-course questionnaire was administered to the original participants to gain additional self-report data.
We followed the university’s guidelines for conducting studies involving human participants. Prior to participation, students were provided with a research notification and a GDPR-compliant privacy notice detailing how their data were to be used in this study. Participating in both studies was fully elective, and participants could leave the study at any time. Participation was based on informed consent, and whether a student participated in the study carried no incentives or penalties. At our university, informed consent is sufficient except when the collected data are categorized as sensitive (race or ethnic origin; political opinions; religious conviction; philosophical beliefs; trade union membership; health information; sexual orientation or behavior; or genetic and biometric data when used to identify a person). Explicit ethics approval is required only when data collection deviates from the informed-consent process, when research requires physical intervention, when the subjects are minors without parental consent, when the research exposes participants to strong stimuli, when it risks causing mental harm, or when it could threaten the safety of participants. As we did not collect sensitive data and none of the other criteria apply, we did not seek a third-party ethics review.
Altogether, $N=47$ students participated in the initial study. None of the students formally withdrew from the study during the course. However, only 29 students also answered the post-course questionnaire, potentially because participation was not mandatory and some participants dropped out of the course. We still included the post-course questionnaire data, as it provided valuable information for explaining student learning behaviors. As such, we used trace log data from all 47 students to investigate the learning behaviors (RQ1 and RQ2) and the questionnaire data from the 29 students to compare self-reports and behavior data (RQ3).
@@ 3.2.1. Data collection
The main data of the study are the trace logs of the participating students. The initial logs were collected as raw action logs which contained timestamps of every action a student performed in the virtual learning environment during the 11 weeks of the course. The raw logs were processed into final trace logs by removing actions irrelevant to the course and by recoding the actions into context-relevant event codes. The specific approach for processing the raw logs and the description of all event codes are provided in Appendix A. The event codes relate to the use of the learning resources available to the students during CS1, such as the coursebook, lectures, supplementary examples, weekly tasks, group sessions, one-on-one sessions, model answers and answer review sessions (see Section 3.1). For the weekly tasks, the separate event codes were given for interacting with the different task types described in Section 3.1 (introductory, basic, core, bonus, guru), as different tasks play a different role in the course. Additionally, the events for interacting with the current week’s tasks and the previous week’s tasks were differentiated, as students could potentially review and revise the previous week’s tasks based on the tasks review sessions and the model answers. Overall, 82 distinct event codes were used to describe student behavior during CS1.
In addition to the trace log data, task completion and performance data was also collected for the duration of the whole course. The data included timestamps of all task submission attempts sent to the virtual learning environment and the correctness of each submission. General performance data, such as weekly task points before the weekly deadline, task points after the deadline (as students could edit their answers based on the model answers), final exam points and final course grade were also collected.
@@ 4. Data analysis pipeline
Figure 2. The pipeline for analyzing trace logs in the present paper. The pipeline consists of five steps, where the outputs of the previous steps are used as inputs for the subsequent steps. Analysis follows Winne and Marzouk (2019)’s characterization of SRL, whereby student’s learning strategy are characterized as transitions of learning tactics.
To answer the research questions, we employ exploratory methods, such as cluster analysis and conventional content analysis, to identify patterns of student behavior and compare them with self-report data. Our aim was not to establish statistically significant generalizations, but to examine the phenomenon through learning analytics supported by qualitative self-reports.
The overall data analysis pipeline is depicted in Figure 2. The pipeline is based mainly on the prior analysis procedures by Matcha et al. (2019a) with modifications to focus more on learning strategies and their association with dropout risk. The pipeline consists of five steps, where the outputs of the previous steps are used as inputs for the subsequent steps. Steps 1–3 reduce the size and dimensionality of the logs by recoding raw events as learning tactics, while steps 4–5 form the main analysis used to answer the research questions. We outline the steps in detail next. The data used and the full analysis code are available online [^2] URL available upon request from the first author..
4.1. Log preprocessing
Before the main analysis, the input trace log is split based on the time distance between events. This starting step is adapted from the prior analysis procedures of Jovanović et al. (2017) and Matcha et al. (2019a) to preprocess logs into student sessions that represent students’ consecutive active interactions with course materials, such as time spent on a single task. First, we grouped all events in the logs by student and sorted them by timestamp. Next, we split the logs for each student: if the time difference between two events was longer than the cut-off threshold, the first event was set as the end of the previous session and the second as the start of a new session. This yields a set of student session logs, which is the step’s output.
A critical parameter for this step is the cut-off threshold, as it affects the granularity and generalizability of the subsequent analysis (Kovanovic et al., 2015). However, there is no set rule for choosing a threshold: Jovanović et al. (2017) opted for 30-minute sessions citing a previous study, while Matcha et al. (2019a) selected 30 minutes based on a review by Kovanovic et al. (2015) noting that it was one of the most common choices. In turn, Zhidkikh et al. (2023a) used a 45-minute cut-off based on the study context and the nature of the tasks. Here, we selected the cut-off based on context: given the number of weekly tasks available to the students, we estimated that a student was allocated 25 minutes for each task on average. As such, we chose 25 minutes as the cut-off threshold in this study.
Before proceeding to the next step, we removed all student sessions that consisted of a single event. This additional filtering step was suggested by Jovanović et al. (2017) to remove potential outliers caused, for instance, by a student leaving the learning environment open and returning to it later or opening the course environment by mistake. Jovanović et al. (2017) also suggested removing sessions whose length falls outside the 95th percentile to focus on more general patterns. However, we opted not to remove long sessions, as our goal was to characterize the full range of students’ learning behaviors, including those who engaged more deeply with the course materials.
4.2. Learning tactic discovery
Once the student sessions are obtained, they are used as input to identify common learning tactics used throughout the course. The core idea of this step is based on the work of Matcha et al. (2019a), whereby learning tactics are identified by applying cluster analysis to student sessions and interpreting the resulting clusters. In addition to identifying learning tactics, this reduces the dimensionality of the data by grouping similar sessions together. The output of this step is a set of clusters of student sessions and their characterizations as learning tactics.
In the original approach by Matcha et al. (2019a), each student session is first converted into a First-Order Markov Model (FOMM) and represented as a transition matrix containing the transition frequencies between consecutive events. These FOMM representations are then clustered using the Expectation Maximization (EM) algorithm, which is useful as it does not require an explicit similarity metric between two student sessions (Aggarwal and Reddy, 2014). A potential concern with using FOMMs relates to the number of events: as the number of events grows, FOMM representations grow quadratically with the number of distinct event types. This can lead to sparsity issues, as each individual session likely contains only a small subset of possible event transitions. Reliable clustering may therefore require a larger number of student sessions. For example, while Zhidkikh et al. (2023a) included 68 sessions with 14 event types, Matcha et al. (2019a) included nearly 5,000 sessions with only 22 event types. Many other studies similarly limit their analysis to roughly a dozen event types (e.g., Villalobos et al., 2022; Fincham et al., 2019; Fan et al., 2021a), a number that rarely matches the range of actions a student can actually perform. In this study, we opted for 82 event codes to obtain more detailed logs (see Appendix A). This presents a possible concern for FOMM-based clustering, as the number of student sessions could be much lower than the number of potential event transitions. Further, a potential limitation of EM is its computational complexity and scalability to large datasets (see e.g., Karimi et al., 2019).
Given these concerns, we considered an alternative approach for identifying learning tactics. Instead of clustering FOMM representations, we represented each student session as an event frequency vector $(\text{fE}_{1},\text{fE}_{2},\ldots,\text{fE}_{82})$, where $\text{fE}_{i}$ indicates the relative frequency of an event type in the session. While this discards information about event transitions, this representation substantially reduces the dimensionality and sparsity of the student sessions. It also allowed us to employ a $k$-means-based algorithm for clustering the student sessions. Because each session vector represents a discrete distribution of event frequencies, we used the $k$-medoids algorithm (Schubert and Lenssen, 2022), which allows applying the $\ell_{1}$ distance metric that is well-suited to such distributions. Once the clusters were obtained, they were reviewed and interpreted by the first author to describe the learning tactics.
4.3. Learning strategy formulation
Using the identified learning tactics as input, the original student sessions are labeled based on the cluster to which they were assigned. We also extracted the timestamp of the first event of each student session and determined the course week in which the session occurred, yielding a trace log of learning tactics that is essentially a higher-level view of the original trace log. Since we defined learning strategies as meaningful sequences of learning tactics (see Section 2.1), we repeated the splitting process of Step 1 on the new learning tactics trace log. However, instead of using a cut-off threshold, we split the trace logs into weekly sublogs. This was done because the course was structured into weekly modules and students were likely to adjust their strategies weekly in response to the weekly context and prior performance (e.g., Pintrich, 2000). This additional splitting yielded per-student weekly learning tactic logs that capture how each student adjusted their tactics throughout the week. The output of this step is thus a set of weekly learning tactic logs, i.e., learning strategies (see Section 2.1).
4.4. Learning strategy characterization
This step aims to characterize the weekly learning strategies and group them into high- and low-risk strategies based on their possible association with dropout. The output of the step is a set of descriptions of the learning strategy types students employed in CS1 and their association with dropout risk, addressing RQ1 and RQ2.
Assessing dropout risk for weekly strategies can be challenging. For one, true dropout can be known only after the course ends. Further, all students adjust their learning strategies in response to the environment and prior experiences (e.g., Pintrich, 2000), which means that even a student who eventually drops out will likely have employed some of the same strategies as students who persist. Similar weekly strategies may therefore be associated with both dropout and persistence. Finally, while manual evaluation of learning strategies is possible (see e.g., Liao et al., 2019), it is not scalable to a large number of students or strategies.
To address the issue of labeling strategies as high- or low-risk, we used a pre-trained dropout risk prediction model based on task completion data. Specifically, we used a pre-trained logistic regression models for classifying dropout developed by Zhidkikh et al. (2024b). These models were trained on a large dataset in a similar CS1 context and were demonstrated to be transferable to other similar course contexts. Importantly, (Zhidkikh et al., 2024b) include models for predicting dropout using only the task performance data available during a specific week of the course. In other words, these models can predict a student’s dropout risk during a given week. The models take as input the weekly task performance data described in Section 3.2.1 and output a dropout prediction value $p_{\text{drop}}$ that ranges from 0 (likely to persist) to 1 (likely to drop out).
In this step, we computed the weekly dropout predictions for each student with the models and used them as a measure of dropout risk associated with the week’s learning strategy. We then labeled a student’s weekly learning strategy as high risk if the model’s prediction for that week was greater than 0.5 and as low risk otherwise. In other words, we classify a strategy as low or high risk by taking the most probable outcome predicted by the model—this is a common approach to deciding the class based on a probability (Hastie et al., 2009, p. 21), and the method originally used to evaluate these models in (Zhidkikh et al., 2024b). Overall, using a prediction model as a measure of dropout risk allowed us to group the weekly learning strategies into high and low dropout risk strategies while accounting for the challenges outlined above.
Once the weekly learning strategies were grouped into high- and low-risk, we clustered the strategies within each group. Here, we proceeded with the same approach Matcha et al. (2019a) use to cluster learning tactics. Each weekly strategy—described as a log of learning tactics—was converted into a FOMM and then clustered using the EM algorithm. We opted for this approach here because reducing the original trace logs into learning tactics significantly reduced the number of potential transitions between tactics, making the EM algorithm suitable for the task. This yielded two groups of clusters: clusters of low dropout risk strategies and clusters of high dropout risk strategies.
To interpret and characterize the clusters as learning strategy types, we produced aiding visualizations. For each cluster, these visualizations displayed the median dropout probability of the associated weekly strategies, a frequency diagram of the tactics present in the cluster, and a visual representation of the most relevant learning tactic transitions across all learning strategies in the cluster. For the latter, we generated a heuristic net (Weijters et al., 2006), a process mining algorithm used to visualize patterns in event logs. A core benefit of heuristic nets is the ability to filter out less frequent transitions, simplifying the visualization of event logs (Weijters et al., 2006). For each learning strategy cluster, we generated a heuristic net from the union of all learning tactic transitions present in the cluster. An example of a heuristic net is shown in Figure 3(a). In addition to heuristic nets, we also generated classic directly-follows graphs for each weekly strategy to represent the exact transitions between learning tactics (Berti et al., 2023). Together, heuristic nets and directly-follows graphs served as a graphical aid in characterizing the learning strategy clusters.
To further aid interpretation, we also examined how the dropout prediction model classified the weekly strategies in each cluster. For this, we used SHAP analysis (Lundberg and Lee, 2017), a general method for explaining the predictions of machine learning models. For each prediction, SHAP analysis assigns an importance value to each feature, indicating how much and in which direction a feature contributed to the dropout prediction. Here, we used SHAP analysis to compute how weekly metrics related to task performance contributed to the weekly dropout prediction. We viewed the feature importances using a waterfall plot (see Figure 3(b) for an example) for each learning strategy in the cluster. Waterfall plots provided a quick overview of how a dropout prediction for each student-week was formulated by the prediction model, particularly in terms of which features contributed most. For example, in Figure 3(b) the strongest explanators of the dropout prediction were the number of correct answers in the previous week and the number of submissions during the current course week. These explanations allowed us to more effectively characterize students’ learning strategies without relying solely on the learning tactics.
(a) Heuristic Net computed and optimized from all strategies in a single learning strategy cluster. The net displays the most relevant transitions between learning tactics from the start of the week (green node) to the end of the week (red node). The net is an aggregation of all learning strategies in a learning strategy cluster.
(b) SHAP waterfall plot for the dropout prediction associated with one of the learning strategies (identified via student identifier and week number) in a learning strategy cluster. The waterfall plot displays dropout probability used to classify a specific learning strategy and how each of the feature contributed to the prediction. Negative values indicate features that explain towards dropout and positive values indicate features that explain towards persistence.
Figure 3. Example of graphs used to identify and characterize learning strategy types from clusters of learning strategies.
The interpretation was performed by the first two authors through two hours of joint work, during which the clusters were reviewed and the interpretations agreed upon using the visualizations described above. In unclear cases, the authors also referred to students’ self-reports to better understand the context of a specific student’s strategy. The resulting strategy types were then used to describe the learning behaviors associated with low dropout risk (RQ1) and with high dropout risk (RQ2).
4.5. Student characterization
The final step of the analysis aims to characterize students’ general activity throughout the course in order to compare this activity with self-reports (RQ3) and to provide further information for characterizing the learning behaviors in RQ1 and RQ2. This step follows the analysis used by Jovanović et al. (2017) and Zhidkikh et al. (2023a). First, the weekly learning strategies identified in Step 3 and the learning strategy types from Step 4 are taken as input. Each student is then assigned a learning profile represented as a frequency vector $(\text{fStrat}_{1},\text{fStrat}_{2},\ldots,\text{fStrat}_{N})$, where $\text{fStrat}_{i}$ represents the frequency of a student’s weekly strategies corresponding to a specific learning strategy type and $N$ is the number of strategy types identified earlier. These profiles capture students’ overall activity and their weekly strategy adjustments during the course. The use of frequencies simplified the profiles and enhanced the robustness of the analysis. Following the analysis described in (Jovanović et al., 2017; Zhidkikh et al., 2023a), we used agglomerative clustering with the $\ell_{1}$ metric to cluster students based on their learning profiles. Agglomerative clustering allows drawing a dendrogram that visualizes the hierarchical structure of the clusters, supporting the selection of a suitable cluster count and the detection of potential outliers. These clusters were then interpreted and characterized as student learning profile types that capture different patterns of learning strategy use throughout the course.
@@ 4.5. Student characterization
To aid in interpreting the student learning profile types, the open-ended questionnaire answers were preprocessed at this stage. The last two authors of this study individually coded the questionnaire answers into themes. The coding was not guided but emerged from the answers (cf. conventional content analysis in Hsieh and Shannon, 2005). The two authors also recorded whether each student expressed a positive, neutral or negative opinion of their own learning behavior during the course. After individual coding, the two authors convened in a joint two-hour session to construct shared interpretations across the entire dataset, resolving discrepancies rather than taking votes. The resulting codes were thus jointly decided between the last two authors, not merely reconciled.
In addition to the self-report codes, we used performance-related metrics for each cluster, such as the dropout probability across all student sessions in each cluster and the final course grades of the students in the cluster. These values, along with the self-reported opinions about learning behavior, were used to compare the clusters and characterize the student learning profile types. To support this mostly qualitative comparison, we employed a statistical test to gauge the differences between the three metrics across the clusters. Here, we sought a test that can accommodate a small number of data points and is robust enough to provide a numerical measure of difference between clusters as an aid. As such, we opted for the Brunner-Munzel test (Neubert and Brunner, 2007), a robust non-parametric alternative to the Mann-Whitney U test that is suitable for small numbers of observations without requiring assumptions about their distribution (Wilcox, 2017; Karch, 2021). Like the Mann-Whitney U test, the Brunner-Munzel test assesses stochastic equality—that is, whether the probability of an observation from one distribution being greater than an observation from another distribution is not random (Karch, 2021). Importantly, the Brunner-Munzel test statistic, the stochastic superiority $\hat{p}^{\prime\prime}$, can be interpreted as the probability that random variable $X$ is lower than random variable $Y$ when ties (i.e., observations of equal value) are split evenly, making it a useful effect size measure (Karch, 2021). We therefore use the Brunner-Munzel test and its effect size $\hat{p}^{\prime\prime}$ to provide a numerical comparison between clusters that aids interpretation.
Overall, using the self-report codes, the self-reports, predicted dropout and course grades as a guide, the first two authors iteratively interpreted the resulting clusters in up to three hours of joint work. These characterizations of student learning profile types were then used to compare self-reports with the behaviors obtained from the trace logs (RQ3).